### v2

# Context Is Not Comprehension

Abstract
The dominant evaluation of Large Language Models has centered on their ability to surface explicit facts from increasingly vast contexts. While today’s best models demonstrate near-perfect recall on these tasks, this apparent success masks a fundamental failure in multi-step computation when information is embedded in a narrative. We introduce Verbose ListOps (VLO), a novel benchmark designed to isolate this failure. VLO programmatically weaves deterministic, nested computations into coherent stories, forcing models to track and update internal state rather than simply locate explicit values. Our experiments show that leading LLMs, capable of solving the raw ListOps equations with near-perfect accuracy, collapse in performance on VLO at just 10k tokens. The VLO framework is extensible to any verifiable reasoning task, providing a critical tool to move beyond simply expanding context windows and begin building models with the robust, stateful comprehension required for complex knowledge work.
1 Introduction
Despite boasting million-token context windows, today’s leading Large Language Models (LLMs) can fail at reasoning tasks a human finds trivial: tracking a multi-step argument buried within a distracting narrative. This ability—to filter irrelevant information and track intermediate conclusions for later synthesis—is the foundation of high-value knowledge work, from a lawyer interpreting intepreting clauses to a sales manager inferring customer intent from chatty transcripts. Yet this core skill remains a fundamental limitation as existing benchmarks are ill-equipped to measure it. To bridge this gap, we introduce Verbose ListOps (VLO), a novel benchmark that isolates this ability in a controlled setting, challenging models to reason, not just locate information.
Figure 1: VLO embeds ListOps problems in a narrative, forcing LLMs to find and solve each subproblem, track intermediate states, and ignore distractors for the final answer, emulating complex human text synthesis. Takeaway: VLO reveals LLMs needs more than just long-context processing—they must compute within distracting narratives where intermediate results remain implicit.
VLO is specifically designed to address the limitations of prior evaluation methods. Where benchmarks for sequence comprehension (An et al., 2023; Zhang et al., 2024) or needle-in-a-haystack factual recall (Kamradt, 2023; Li et al., 2024b) test factual extraction, VLO embeds deterministic, nested ListOps computations (Nangia and Bowman, 2018; Tay et al., 2021) inside lengthy, coherent stories. Where synthetic datasets can leak intermediate solutions that models merely recall (Fodor, 2025), VLO enforces a strict protocol that withholds every intermediate result, forcing models to compute and maintain values in ’working memory’. And while human-annotated datasets offer realism, they are too coarse-grained to disentangle narrative interference from reasoning difficulty (Wang et al., 2024; Shaham et al., 2022; Bowman and Dahl, 2021); VLO’s agentic generation provides orthogonal controls for both. This setup probes an LLM’s ability to maintain computational ’state’ and follow algorithmic logic under narrative camouflage—coherent, semantically-related distractors that obscure a task—prerequisites for text understanding comparable to human performance.
Benchmark Primary Reasoning Task Nested Reasoning Chains Distraction Tunable Reasoning Difficulty Scalable Context Deterministic Answer Realistic Tasks Generation Method
Verbose ListOps (Ours) Algorithmic (ListOps) embedded in narrative Coherent, semantically relevant narrative Mathematically deterministic Agentic
LongReason (2025) General QA (Comprehension, Inference, Maths) $\times$ Irrelevant passages embedded around relevant context $\approx$ Multiple-choice (via reasoning) Context expansion
Needle-in-Haystack Type Factual Recall $\times$ Vast irrelevant corpus $\times$ Exact match $\times$ Target insertion
Other Synthetic (e.g. RULER 2024.) Code/Instruction Following $\times$ Structured, less narrative $\times$ Task-specific $\times$ Template Based
Human Annotated Long-Context QA General QA $\times$ Natural document structure $\times$ $\times$ Human-judged Human annotation
Table 1: Comparison of Verbose ListOps with other long-context benchmarks. Using an agentic generation process, Verbose ListOps offers both controllable context lengths and reasoning difficulty.
In this paper, we deliver three key contributions:
  1. 1.
    A benchmark for narrative-embedded deterministic reasoning. We weave ListOps computations into coherent fictional stories and withhold the intermediate results of all the nested problems, forcing models to compute, store, and recall values internally.
  2. 2.
    Orthogonal control of context length and reasoning complexity. We provide independent parameters for narrative length and reasoning complexity, enabling systemic exploration of scaling behavior and failure modes.
  3. 3.
    An extensible generalizable generation framework. We open-source a fully programmatic agentic pipeline where "author" and "critic" LLMs collaborate to embed any deterministically verifiable task; from numerical operations to symbolic and logical reasoning, into a coherent narrative. This methodology yields reliable, scalable data and provides a foundation for testing a wide range of complex reasoning beyond the scope of this initial work.
Our experiments show state-of-the-art LLMs, despite solving ListOps with ease, on VLO, suffer a severe ($\approx$50%+) drop in accuracy at just $\approx$10k-token contexts. The result exposes an under-tested limitation: LLMs struggle to synthesize multi-step conclusions amongst narrative distraction. VLO offers a rigorous, open-source tool for diagnosing (and improving) this foundational capability, seeking to move discussion beyond simply enlarging context windows.
2 Related Work
The expansion of LLM context windows has catalyzed research into their performance on extensive textual inputs, leading to a diverse landscape of evaluation benchmarks.
2.1 Long-Context Benchmarks
Early long-context evaluations typically adapted standard NLP benchmarks (Shaham et al., 2023; An et al., 2023; Bai et al., 2024). These early benchmarks often had contexts shorter than the maximum capabilities of current LLMs (Gemini Team, 2024; OpenAI, 2025a; Anthropic, 2025) and inadequately differentiated complex reasoning tasks from simpler simpler tasks based on factual recall tasks or failed to analyze the impact of distractors sufficiently (Muhlgay et al., 2023).
Recent synthetic benchmarks have addressed some of these limitations by allowing greater control over context length. The "needle-in-a-haystack" (NIAH) paradigm, exemplified by NeedleBench (Li et al., 2024b), specifically evaluates an LLM’s recall capability within extensive irrelevant contexts, focusing on fact extraction rather than complex multi-step reasoning. Other synthetic benchmarks, such as RULER (Hsieh et al., 2024), test specific reasoning capabilities like variable tracking and multi-hop information extraction. However, their artificial scenarios may not adequately replicate the challenges posed by naturalistic narratives with coherent distractors (Haller et al., 2024).
Benchmarks such as InfiniteBench (Zhang et al., 2024) or those utilizing extensive document curation like LooGLE (Li et al., 2024a) often rely heavily on human annotation or semi-automated methods. While realistic, these approaches are labor-intensive and limit scalability and precise control over task variables such as context length and complexity, underscoring the need for standardized, synthetic evaluation frameworks (Valmeekam et al., 2023; Muhlgay et al., 2023).
LongReason (Ling et al., 2025) significantly advances synthetic benchmarks by offering variety—reading comprehension, logical inference, and mathematical reasoning tasks are expanded into longer, distractor-rich texts. It challenges models to aggregate and reason over scattered information. In contrast, the VLO benchmark presents a distinct challenge by embedding nested algorithmic ListOps problems within coherent narratives, uniquely emphasizing internal computation and state management. This approach differs fundamentally from LongReason, which centers on aggregating explicitly presented clues rather than tracking values that must be computed from prior steps.
2.2 The Emergence of Reasoning alongside Long Context
As strong large-context capabilities became the norm, demand grew for more complex reasoning within those extended windows, fueling the development of "reasoning" models and benchmarks. Surveys of long-context benchmarks show LLMs recall facts across tens of thousands of tokens but falter at multi-step inference over extended inputs. Liu et al. (2024) find ultra-long transformers (100k+ tokens) struggle to integrate dispersed information as context grows. Hsieh et al. (2024) report similar declines on RULER tasks, highlighting that wider windows alone don’t guarantee accurate intermediate-state tracking. Shi et al. (2023) term this "attention dilution," where relevant facts become difficult to aggregate in long contexts. Consequently, current research now emphasizes explicit planning, modular computation, or memory mechanisms.
Prompting-Based Reasoning Techniques.
The first widely adopted method to elicit stepwise reasoning from LLMs was Chain-of-Thought (CoT) prompting, instructing models to think step-by-step or via in-context examples, enhancing arithmetic and logic performance without fine-tuning (Wei et al., 2022). However, CoT becomes brittle with deep nesting or backtracking, causing cascading errors. Yao et al. (2023) propose Tree-of-Thoughts, which explores multiple reasoning branches in parallel, uses verification to prune incorrect paths, and allows backtracking. This search-based method helps models recover from missteps and leverage dispersed information more effectively.
Architectures for Explicit Reasoning.
Beyond prompt engineering, recent architectures directly embed planning and state-tracking. Large Concept Models Meta LCM Team et al. (2024) process higher-order semantic units rather than text tokens, tracking each reasoning step as a ’concept’. Similarly, Mondal et al. (2024) introduce a prefrontal cortex–inspired Modular Agentic Planner (MAP), where a central LLM orchestrates modules proposing, evaluating, and refining subgoals via a shared scratchpad to track intermediate states. Outperforming single-pass transformers on classical benchmarks (e.g., Tower of Hanoi), MAP highlights benefits to decoupling planning from execution.
Memory-Augmented and Retrieval-Augmented Reasoning.
Architectures with explicit memory modules show promise over long-context reasoning. Recurrent Memory Transformers (RMTs) carry summaries of past hidden states across segments to avoid reprocessing. In BABI-Long evaluations, RMTs outperform standard long-attention models by integrating clues over 128k tokens (Wang et al., 2024). Retrieval-augmented generation (RAG) complements these: by fetching relevant passages, LLMs ground inferences (e.g., An et al. (2023)). However, Hengle et al. (2025) show RAG pipelines—despite high retrieval accuracy—often fail when chained reasoning steps are needed in extended contexts. Likewise, the LaRA benchmark (2025) finds long-context LLMs outperform RAG on multi-step arithmetic problems as document length scales (Li et al., 2025).
3 Verbose ListOps (VLO): Specification and Construction
This section formalises the Verbose ListOps (VLO) task, details its programmatic construction pipeline, and presents the controllable parameters used to stress-test long-context language models.
3.1 Formal Task Definition
Given a narrative $X\in\mathbb{T}_{10\mathrm{k}\text{-token}}$ and an implicit ListOps abstract syntax tree (AST) $T$ with leaves (atomic integers) and internal nodes (operators), the original ListOps task (Nangia and Bowman, 2018) requires models to evaluate $T$ by performing its specified operations (e.g., sums, minimums, medians) on numerical inputs to yield a deterministic result. For example, if
$T=\textsc{max}(\textsc{sum}(2,1),\,4),$
then one computes $\textsc{sum}(2,1)=3$ and subsequently $\textsc{max}(3,4)=4$.
VLO embeds each nested ListOps operation in a narrative, describing each node in $T$ via post-order traversal. Placeholders $\{a_{i}\}$ (which we call ’narrative anchors’) refer to intermediate values without explicitely stating the results.
Formally, the VLO task is:
$f\colon X\longrightarrow v(T),$
with intermediate values appearing only as anchors ${a_{i}}$. Success is $f(X)$ matching the ground-truth integer.
Running example
Consider the above ListOps problem again. The first operation is
$\textsc{sum}(2,1)\;\to\;a_{1},\quad a_{1}=3$
This intermediate value (3) is tracked by anchor $a_{1}$. Example narrative segment:
"In the bustling spaceport of Xylos, Chief Engineer Anya cataloged incoming parts. Her logs showed one crate of cryo-cells arrived on the morning freighter, and later, two additional crates were offloaded from a fast courier. She processed these figures, which she labeled as ’Daily Cell Intake’ (anchor $a_{1}$) in preperation for her presentation at Xylos’ annual Meteor Shower Festival.
Notice ’Daily Cell Intake’ represents $a_{1}$ (the first operation’s result), done so the number ‘3’ (result of the sum) never appears.
The second operation is
$\textsc{max}(a_{1},4)\;\to\;a_{2},\quad a_{2}=4$
…yielding the final result $v(T)$. The narrative then compares ’Daily Cell Intake’ ($a_{1}$) with four, names the larger as anchor $a_{2}$, and omits that $a_{1}=3$. This preserves the no-numeric-leakage constraint and ensures the model’s output equals $v(T)$.
3.2 Design Constraints
VLO enforces three orthogonal constraints that distinguish it from prior long-context benchmarks:
C1 No numeric leakage.
Intermediate results are never stated; they may only be referenced via anchors (e.g., “the daily cell intake”). Explicit numerals for non-leaf nodes are prohibited.
C2 Narrative camouflage.
LLMs weave operator descriptions into a coherent story that contains LLM-generated, semantically related computation-irrelevant content (distractors), mirroring real-world documents where crucial information is buried among related context.
C3 Parametric control.
ListOps complexity (operator set, tree depth $d$, branching factor $b$) is adjustable, enabling ablations; increasing $d$ or $b$ adds anchors and narrative density.
3.3 Generation Pipeline
Figure 2 outlines how the VLO construction pipeline ensures adherence to constraints C1–C3 through:
  1. 1.
    AST Sampling (Programmatic). Sample a ListOps AST $T$ of depth $d$ and branching factor $b$, and compute its root value $v(T)$ by deterministic post-order evaluation.
  2. 2.
    Iterative Operator Narrative Construction. For each ListOps node, build a narrative segment via:
    • (a)
      Author Agent (LLM). The Author LLM drafts a segment for the current operator $\omega\in\mathcal{O}$ and its child anchors $a_{i}$ (if applicable), weaving it into the story (C2) while avoiding numeric mention of the current result (C1) and respecting parameters (C3).
    • (b)
      Critic Agent (LLM). The Critic LLM reviews the draft, checking:
      • No numeric digits or words for the current operation’s result appear (enforces C1).
      • Correct anchor use (for child inputs of prior operations) and atomic inputs.
      • Narrative coherence (ensures C2).
      It proposes edits if it finds violations, and the Author revises until all checks pass.
    • (c)
      Static Validation (Programmatic). A script scans the finalized segment to verify:
      • Absence of any digit or word numbers for the current operation’s result.
      • Presence of anchor tokens (if expected) or numerical atomic inputs.
      • No unexpected placeholders or formatting.
      Any failure triggers a re-generation of that segment.
  3. 3.
    Holistic Narrative Validation (LLM). After concatenating all segments into the full narrative, validator.py orchestrates a review of the entire sample:
    • Comprehensive LLM validation: A larger state-of-the-art Validator LLM, aided by extensive prompting and few-shot examples, reads the entire story. It performs a step-by-step evaluation of the narrative against the original AST $T$ to detect subtle numeric leaks (e.g., spelled-out numbers for intermediate results), verify correct representation of operations and inputs (including conceptual anchors), check for missing or duplicate operators, ensure overall coherence (C2), and confirm that its own evaluation of the original AST $T$ matches the provided ground truth value $v(T)$.
    Violations identified by the Validator LLM cause the sample to be discarded and resampled.
Figure 2: The Verbose ListOps (VLO) agentic generation and validation pipeline, employing Author, Critic, and Validator LLMs alongside programmatic checks to ensure sample validity.
3.4 Controllable Parameters
VLO’s generation pipeline allows fine-grained difficulty control via Config. Key parameters include:
  • Narrative length: The target length of a sample.
  • Tree complexity: Parameters such as tree depth $d$ (MAX_OPS) and operator branching factor $b$ (arity, controlled by MIN_ARITY and MAX_BRANCH) are tunable. For the specific settings used to generate the dataset for this paper, please refer to Section 3.6 and Appendix C. Larger $d$ or $b$ generally increases the number of anchors and thus the narrative density.
3.5 Evaluation Protocol
Models receive the full narrative $X$ (with anchors $\{a_{i}\}$) and must output an integer. We report exact-match accuracy with 95% Wilson confidence intervals, as in recent long-context benchmarks  (An et al., 2023; Zhang et al., 2024). evaluator.py conducts the evaluation (and equation_llm_evaluator.py for standard ListOps). Chain-of-thought or external tool use is prohibited to isolate internal computation. All code, generation logs, and datasets are open-source.
3.6 Experimental Setup
  • Models. We test Gemini 2.5 Pro and Flash, OpenAI o4-Mini and GPT-4.1, Grok 3-Mini, Claude 3.7 Sonnet, DeepSeek R1 and V3, Qwen-3 235B, and Llama-4 Maverick.
  • VLO instances. 1,000 VLO-10k samples were generated at a cost of \$$\approx\$1,500$ USD. The embedded ListOps problems within these instances used the following parameters: max 8 operations per AST, max branching factor of 8, min operator arity of 4, and atomic integer values from range [1, 30]. Full hyperparameter details are available in Appendix C.
  • Standard ListOps baseline. For each VLO sample, we construct the bare ListOps expression, providing an algorithmic baseline of identical difficulty without narrative distractors.
  • Evaluation inference config. All models had $temperature=0.01$ and $top\_p=0.95$ and were accessed via OpenRouter. Total evaluation incurred a cost of \$$\approx\$500$ USD.
3.7 Results
The table below reports exact-match accuracy on both Verbose ListOps and its corresponding bare ListOps expressions alongside performance on the OpenAI MRCR NIAH benchmark. Whilst nearly every ’thinking’ model achieves perfect or near-perfect accuracy on bare ListOps, performance collapses under Verbose ListOps at just 10-k tokens.
Model ListOps VLO-10k OAI MRCR-8k 8-Pin 95% CI (ListOps) 95% CI (VLO)
Closed-source models
Gemini 2.5 Pro 100.0 55.3 86.0 99.6–100.0 52.2–58.4
Gemini 2.5 Flash Thinking 100.0 49.1 63.3 99.6–100.0 46.0–52.2
o4 Mini High 100.0 48.1 63.6 99.6–100.0 45.0–51.2
Grok 3 Mini High 81.7 47.2 51.6 79.2–84.0 44.1–50.3
Claude 3.7 Sonnet High 58.4 40.0 N/A 55.3–61.4 37.0–43.1
GPT-4.1 38.6 32.2 29.6 35.6–41.7 29.3–35.2
Open-source models
DeepSeek R1 98.4 41.2 N/A 97.4–99.1 38.2–44.3
Qwen-3 235B 53.7 41.5 N/A 50.6–56.8 38.4–44.6
Llama 4 Maverick 47.2 32.0 N/A 44.1–50.3 29.1–35.0
DeepSeek V3 0324 93.7 25.1 N/A 92.0–95.1 22.4–28.0
Table 2: Accuracy on VLO-10k versus its corresponding bare ListOps expressions, based on evaluation of 1,000 samples per model. Of particular interest is DeepSeek V3’s plunge in performance from 93.7% to 25.1%. OpenAI MRCR results (95% CI) sourced from ContextArena (Uzar, 2025).
3.8 Analysis and Limitations
Narrative Camouflage, ’Cognitive Control’, and Architectural Divergence
The VLO benchmark’s core challenge is its use of narrative camouflage to obscure a nested deterministic task, creating a information processing burden that exposes fundamental architectural differences between models. The results in Table 2 reveal two critical takeaways. First, models with explicit reasoning scaffolds (e.g., Gemini 2.5, o4-Mini) perform better on VLO than those without. Second, the stark contrast between DeepSeek-V3’s near-perfect score on raw ListOps and its collapse on VLO demonstrates while scratchpad-like mechanisms are not required for nested reasoning, they are crucial for filtering noise and protecting the reasoning process from distractors.
This architectural divergence explains the performance trends. DeepSeek-V3’s design, while highly efficient, creates specific vulnerabilities to VLO’s challenge. Its architecture is optimized for performance on standard benchmarks through two key strategies that, we hypothesize, compromise its task resiliance on VLO. First, its auxiliary-loss-free load balancing for its Mixture-of-Experts (MoE) architecture encourages over-specialisation, where the gating network routes inputs to a "narrative" expert that fails to pattern-match the embedded logic. Second, its use of Multi-Token Prediction (MTP) encourages the model to "pre-plan" its output based on narrative flow, reinforcing the very heuristic processing that VLO penalizes (DeepSeek-AI et al., 2025).
Crucially, DeepSeek-V3’s technical report clarifies that its math and coding reasoning capabilities stem not from emergence but from distillation from DeepSeek-R1, which has long-Chain-of-Thought capabilities. This process "notably improves its reasoning performance" by adopting R1’s "verification and reflection patterns" into V3 (DeepSeek-AI et al., 2025). Therefore, V3’s reasoning is a specialized, distilled skill rather than an inherent, flexible process—explaining its brittleness: these patterns excel on structured tasks like raw ListOps but falter under VLO’s narrative camouflage, which demands a more robust, first-principles reasoning scaffold.
In contrast, Gemini’s architecture, founded on a highly efficient sparse Mixture-of-Experts (MoE) architecture (Gemini Team, 2024), appears to incorporate more robust mechanisms for maintaining task focus amidst distraction. An analysis of Gemini 2.5’s ’thinking’ on VLO suggests its MoE implementation may avoid this hyper-specialization trap. When processing a VLO problem, it behaves as if it first decomposes the task, managing the narrative, operands, and operators as distinct variables. Its MoE routing is then conditioned on this sub-task, and behaves in a way that seems to dynamically shift from activating semantic experts for the narrative to logic-and-reasoning experts for the calculation, in what can be described as dynamic context-aware routing. This ability to parse and apply a formal, rule-based system provided entirely in-context is the same fundamental capability demonstrated in the Gemini 1.5 Technical Report, where the model learns to translate Kalamang—a language with <200 speakers—by processing an in-prompt 500-page grammar book and dictionary (Gemini Team, 2024).
Viewing this delta as a proxy for maintaining appropriate attention over long contexts, this architectural split can explain the shrinking performance delta between MRCR and VLO for less capable ’thinking’ models. Less capable models have uniformly weak heuristic processing—not sophisticated enough to be hijacked by narrative—and too weak algorithmic execution, yielding two low, closely clustered scores. Conversely, Gemini 2.5’s architecture explains its large delta: its heuristic system excels at MRCR but generates interference that its internal reasoning scaffold struggles to overcome on VLO, a known LLM issue (Shi et al., 2023; Vishwanath et al., 2025). This aligns with findings that LLMs struggle to shift from intuitive, pattern-matching reasoning to deliberate, step-by-step processing for unfamiliar or complex structures (Mirzadeh et al., 2024). Thus, the performance gap between recall and narrative reasoning powerfully proxies this architectural conflict.
Limitations and Future Work
Using a fixed $\approx$10k-token narrative exploits the strong recall abilities of SOTA LLMs at this context size (Uzar, 2025), allowing VLO to isolate reasoning deficits from failures in factual recall. The degradation observed (Table 2) at a context length where recall is robust underscores VLO’s focus on narrative-embedded computational challenges. Future studies should vary context length more broadly to explore these dynamics.
Further considerations and avenues for future research include:
  • Generator Model Bias: VLO-10k was generated with Gemini 2.5.111During development we found only Gemini 2.5 and GPT-4.5 could reliably generate VLO, with GPT-4.5 pricing being cost-prohibitive (\$75USD/1M input tokens, \$150USD/1M output tokens (OpenAI, 2025b)). This introduces potential generator model bias, where evaluated models from the same family may exhibit inflated performance due to stylistic or implicit knowledge alignment. Future work should investigate this and explore mitigation strategies, such as employing a more diverse set of future generator models or incorporating bias-neutralization techniques (Yuan et al., 2025).
  • Chain-of-Thought (CoT) Prohibition: This evaluation prohibits external CoT prompting to isolate core computational reasoning as today’s ’thinking’ models have internal CoT-like functions. Given DeepSeek v3’s notable degradation on VLO, a future evaluation allowing external CoT or advanced prompting (e.g., Tree-of-Thoughts Yao et al. (2023)) could clarify if explicit scratchpad prompting restores reasoning or reveals inherent architectural limits.
  • Depth of Failure Mode Analysis: While Table 2 compellingly demonstrates significant performance drops, the current analysis remains primarily quantitative. A granular error analysis would help explain why models fail on VLO tasks—whether due to narrative parsing errors, incorrect ListOps computations, or challenges in tracking internal states amid distractors. Investigating error cascading or differences by operator type, tree complexity, or narrative structure could inform the development of more robust reasoning architectures.
  • Scope and Generalizability of ListOps: Currently, VLO uses ListOps to evaluate algorithmic execution and state-tracking within narratives. This initial focus does not capture the full range of real-world narrative reasoning, which involves ambiguity, implicit knowledge, and logical frameworks like abductive, inductive (Sheng et al., 2025; Bowen et al., 2024), or defeasible reasoning (Ren et al., 2024; Leidinger et al., 2024). Fortunately, the VLO generation pipeline is highly extensible, and can support symbolic non-numeric problems, enabling future variants to test reasoning such as abducting causes, inducing general rules, or handling defeasible updates—providing a more comprehensive LLM evaluation. See Appendix A for details on extending VLO to these other reasoning types.
Key takeaway.
Sustaining multi-step computation amidst noise requires architectures that explicitly model reasoning steps. Parameter scale or context window size alone cannot bridge this weakness.
4 Discussion
VLO was initially developed to benchmark automated predictive signal extraction from distributed narratives (e.g., assessing prospect ’consequence awareness’ from sales communications). Results here highlight a critical gap in automating sophisticated analytics: failures observed in VLO mirror real-world issues, such how a lawyer might lose track of interacting clauses or salespeople misinterpreting intent due to narrative complexity. State-of-the-art LLMs excel at fact recall but falter when reasoning with extracted facts, underscoring that improved context length alone is insufficient. VLO thus provides an essential testbed for emerging architectures capable of addressing these reasoning challenges, including those with explicit planning (e.g., PFC-inspired systems (Mondal et al., 2024)) and advanced prompting strategies like Yao et al. (2023)’s Tree-of-Thoughts.
Concept-oriented Large Concept Models (LCMs),222’LCMs’ here do not refer to Latent Consistency Models (Luo et al., 2023) used for image synthesis. which operate on higher-level semantic units rather than individual tokens (Meta LCM Team et al., 2024) offer a promising direction. For VLO, an LCM treats each reasoning step or conceptual reference to an intermediate result (our "narrative anchors") as distinct conceptual units. This aligns more closely with how humans might track such information and could make the model more resistant to narrative distractors when trying to maintain the integrity of a numerical value associated with a concept. In performing autoregressive prediction in a concept embedding space, rather than a token space, LCMs could facilitate more stable tracking of multi-step computations. Such concept embeddings can be a robust and less diffuse carrier of the numerical value of an intermediate result than distributed token activations.
Finally, VLO’s programmatic framework shows promise in how it agentically embeds deterministic tasks into narratives to force internal computation. Such a method can be extended to any domain requiring implicit criteria encoding (done here via narrative anchors) to create more realistic and challenging synthetic datasets, e.g., LLM-as-Judge, Multi-Hop QA, Code-Based Evaluation, and Long-Context QA/Summarization evaluations. VLO thus not only exposes current limitations but also offers a methodology for developing next-generation LLMs with deeper reasoning.
Broader Impacts
VLO pushes for Large Language Models that can effectively process long, complex narratives. Such advances are stepping stones for LLMs to unearth predictive signals hidden in unstructured text. The upsides are clear: sharper analytical tools for science, more insightful legal review, or more astute financial risk assessment from textual data. Our effort to pinpoint and fix current reasoning flaws is about building more dependable Artificial Intelligence for these demanding roles.
However, creating LLMs skilled at deciphering narratives raises significant dual-use concerns. The capacity to extract predictive signals from text, while beneficial, could be repurposed for high-stakes social monitoring. This creates a direct risk of what is often termed predictive policing, where models could generate pre-emptive, and potentially biased, judgments about individuals based on textual data (European Crime Prevention Network (2022), EUCPN). If an AI claims to "predict" behavior from text, the door opens to profiling and discrimination, creating a climate of surveillance that could chill free expression—a core concern in the literature on surveillance capitalism (Zuboff, 2019).
Further, relying on LLMs for prediction is risky when their interal logic is opaque: unexpected failures increase in likelihood, with these errors being hard to detect or correct (Rane et al., 2024). Hence, high stakes ’automated signal extraction’ gone wrong can deeply affect lives. AI-driven economic shifts likewise redefine professional roles through task-based automation rather than simple job loss, polarizing the labor market: routine analytical tasks may be automated while experts who manage, interpret, and validate these systems become more in demand (Autor et al., 2020).
VLO directly contributes to the development of more scrutable and robust AI. Opaque reasoning can lead to biased outcomes or unexplained failures in critical domains—problems that require tools to pinpoint and falsify specific reasoning paths. The risks of opaque reasoning—from biased predictions to unexplainable failures in safety-critical domains—cannot be mitigated without tools that allow for the falsification of specific reasoning pathways. VLO offers such a tool: a controlled, deterministic envrionment where a model’s failures reveal precise breakdowns in its computation rather than just an accuracy drop. By isolating narrative-embedded reasoning, it delivers a clear diagnostic signal for architectural and algorithmic improvements. This supports the community’s shift from merely scaling capabilities to building trustworthy systems (Stanford Institute for Human-Centered AI (2025), HAI). Responsible AI development demands models that are not only powerful but also interpretable—and VLO helps us understand exactly where and why they fail.
Appendix
Appendix A Extending VLO for Other Reasoning Types / Non-Numberic Reasoning
As VLO is built on a deterministically verifiable Abstract Syntax Tree (AST), this core component can be adapted from numerical operations to symbolic and logical ones. The existing framework can be extended to test abductive, inductive, and defeasible reasoning with non-numeric symbols, while remaining deterministically verifiable.
A.0.1 Adapting the Core: The Abstract Syntax Tree (AST)
The current AST uses operators like SUM, MAX, MIN on integer Atom nodes. This can be generalized:
  • Atoms: The Atom node, which currently holds an integer, can be modified to hold a string representing a non-numeric symbol, a fact, or a concept (e.g., “the ground is wet”, “Tweety is a bird”).
  • Operators: The OpNode can be defined with new operators that represent logical reasoning tasks, such as ABDUCE, INDUCE, or DEFEASIBLE_QUERY.
The crucial step is to define how these new operators are evaluated deterministically in the eval_node function.
A.0.2 Making Logical Reasoning Deterministically Verifiable
The main challenge is making subjective-sounding reasoning tasks verifiable. This is achieved by defining a clear, programmatic evaluation logic for each new operator within a constrained environment.
Abductive Reasoning (Inference to the Best Explanation)
  • Goal: To infer the most likely cause given a set of observations.
  • Deterministic Method: Define a simple, score-based logic. The ABDUCE operator would take a set of observations and a list of potential causes, each with predefined properties. The eval_node function would calculate a score for each cause and select the “best” one based on this score.
  • Example AST:
    (ABDUCE
      (OBSERVATIONS "lights flicker" "strange hum")
      (CAUSES
        (CAUSE "power surge" (simplicity 2) (likelihood 0.8))
        (CAUSE "ghost" (simplicity 8) (likelihood 0.1))
  • eval_node Logic: It would compute score = likelihood / simplicity for each cause and return the name of the cause with the highest score. In this case, “power surge” (0.4) beats “ghost” (0.0125). The ground truth is deterministically “power surge”.
Inductive Reasoning (Generalization)
  • Goal: To form a general rule from specific examples.
  • Deterministic Method: Constrain the space of possible rules. The INDUCE operator would take a list of examples and a predefined set of potential rules. The eval_node function would select the first rule from the set that is consistent with all provided examples.
  • Example AST:
    (INDUCE
      (EXAMPLES ("raven A is black") ("raven B is black"))
      (RULE_CANDIDATES ("all birds are black") ("all ravens are black")
                       ("some ravens are black"))
  • eval_node Logic: It would check each rule candidate. “all birds are black” is not contradicted but is less specific. “all ravens are black” is consistent. It would return “all ravens are black” as the correct, most specific, consistent rule from the given set.
Defeasible Reasoning (Rules with Exceptions)
  • Goal: To reason with rules that can be defeated by new information.
  • Deterministic Method: Implement a simple, priority-based logic system. The DEFEASIBLE_QUERY operator would take a set of facts and an ordered list of rules and exceptions. The eval_node function applies the rules in order, allowing later rules (exceptions) to override earlier ones.
  • Example AST:
    (DEFEASIBLE_QUERY
      (QUERY "Tweety can fly")
      (KNOWLEDGE_BASE
        (FACT "Tweety is a bird")
        (FACT "Tweety is a penguin")
        (RULE "all birds can fly" (priority 1))
        (EXCEPTION "penguins cannot fly" (priority 2))
  • eval_node Logic: It would first conclude “Tweety can fly” from the priority 1 rule. Then, it would process the priority 2 exception, which defeats the initial conclusion. The final, deterministic ground truth is False (i.e., “Tweety cannot fly”).
A.0.3 Extending the Generation and Validation Pipeline
With a deterministic AST in place, the rest of the verbose-listops.py pipeline can be adapted:
  1. 1.
    build_random_ast: This function would be updated to construct these new symbolic and logical ASTs from a set of predefined templates to ensure the generated problems coherence.
  2. 2.
    generate_narrative:
    • Prompts: The prompts would be modified. Instead of asking the LLM to narrate a scene about finding the MAX of a set of numbers, you would ask it to narrate a scene where characters reason about the most likely cause of an event.
    • Narrative Anchors: The concept of “narrative anchors” is even more powerful here. The result of an ABDUCE operation (“power surge”) could be given the anchor “The Prime Theory,” which then becomes a symbolic input for a subsequent reasoning step.
  3. 3.
    Validation (make_number_validator and validator.py): This part requires the most significant rewrite, shifting from numerical validation to symbolic validation.
    • The Goal Remains: The core validation goals are the same: ensure all required inputs are mentioned, the (now symbolic) result is kept implicit, and no extraneous information or conclusions are leaked.
    • New Logic: Instead of extract_numbers_from_text, a function like extract_facts_from_text that uses string matching or regex to verify that the narrative mentions would need be added, for example, “the lights flicker” and “a strange hum”.
    • Implicit Result: The validator would check that the word “power surge” is not explicitly stated as the conclusion, but is only referenced by its anchor (“The Prime Theory”).
A.0.4 Summary
verbose-listops.py’s modular, AST-driven, and agentic-validation architecture provides a robust and sophisticated foundation to extend VLO to test deeper, symbolic reasoning by:
  1. 1.
    Defining the symbolic operators and their deterministic evaluation logic.
  2. 2.
    Creating templates for generating coherent, symbolic ASTs.
  3. 3.
    Rewriting the prompt templates in _generate_narrative_recursive to guide the LLM in narrating these logical problems.
  4. 4.
    Replacing the numerical validation logic with a symbolic/factual validation system that enforces the same core principles of operand presence and result implicitness.
The result would be a novel and powerful benchmark that pushes LLMs beyond numerical computation into the realm of structured, verifiable, narrative-based logical reasoning.
Appendix B Models Evaluated
Table 3 lists the Large Language Models (LLMs) evaluated in this study on the Verbose ListOps benchmark. These models represent a range of state-of-the-art closed-source and open-source offerings available at the time of evaluation (16 May 2025).
Table 3: Large Language Models Evaluated on Verbose ListOps.
Model Name Version Max Input Tok. OpenRouter Identifier
Closed-Source Models
[…]
Open-Source Models
DeepSeek R1 2025/01/20 Varies deepseek-r1
DeepSeek V3 0324 Varies deepseek-chat-v3-0324
[…]
Note: Claimed max input tokens are approximate and subject to change based on provider updates. Identifiers are illustrative examples based on common API/HF naming conventions and may vary. The specific models used for evaluation are as listed in Table 2 of the main paper.
Appendix C Verbose ListOps Generation Hyperparameters
The Verbose ListOps benchmark instances were programmatically generated. Key parameters and settings for the generation process are detailed below. These correspond to the Config dataclass and other settings in the verbose-listops.py script.
C.1 Core ListOps Parameters
  • Maximum Operations (MAX_OPS): 8 (Controls the maximum depth/complexity of the ListOps Abstract Syntax Tree).
  • Maximum Branching Factor (MAX_BRANCH): 8 (Maximum number of children for any operation node).
  • Minimum Arity (MIN_ARITY): 4 (Minimum number of children for any operation node).
  • Atom Value Range (MIN_ATOM_VAL, MAX_ATOM_VAL): 1 to 30.
  • Early Termination Probability (EARLY_TERMINATION_PROBABILITY): 0.0 (Probability of terminating Abstract Syntax Tree branch growth before MAX_OPS is reached).
C.2 Narrative Generation Parameters
  • Narrative Generation Large Language Model (MODEL in verbose-listops.py): google/gemini-2.5-flash-preview:thinking (accessed via OpenRouter API).
  • Target Total Tokens (MAX_TOTAL_TOKENS): 10 0001000010\,00010 000 (For the experiments reported in Table 2). The script can generate longer narratives.
  • Padding Max Token Percentage (PADDING_MAX_TOK_PERCENT): 0.75 (Maximum percentage of the *remaining* token budget (after beats) that can be used for padding).
  • Max Padding Paragraphs per Slot (MAX_PAD_PARAGRAPHS): 30.
  • Use Narrative Anchors (USE_NARRATIVE_ANCHORS): True (Conceptual names for intermediate results).
  • Use Large Language Model for Anchor Naming (USE_LLM_NAMING): True.
  • World Generation Parameters:
    • Min/Max Characters (MIN_WORLD_CHARS, MAX_WORLD_CHARS): 6 to 8.
    • Min/Max Concepts (MIN_WORLD_CONCEPTS, MAX_WORLD_CONCEPTS): 3 to 7.
    • World Generation Temperature (WORLD_GEN_TEMP): 0.9.
  • Beat Generation Temperature (BEAT_GEN_TEMP): 0.5.
  • Creative Narrative Temperature (for Intro/Padding) (CREATIVE_NARRATIVE_TEMP): 0.5.
  • Anchor Generation Temperature (ANCHOR_GEN_TEMP): 0.85.
C.3 Iterative Validation and Retry Parameters
  • Large Language Model Validator Model (LLM_VALIDATOR_MODEL): google/gemini-2.5-flash-preview:thinking (Used in the iterative beat generation loop).
  • Large Language Model Validator Temperature (LLM_VALIDATOR_TEMP): 0.05.
  • Beat Revision Temperature (BEAT_REVISION_TEMP): 0.1.
  • Max Large Language Model Validation Iterations (MAX_LLM_VALIDATION_ITERATIONS): 6 (Internal loop for a single beat).
  • Max Beat Retries (Outer Loop) (MAX_BEAT_RETRIES): 5.
  • Max Padding Retries (MAX_PAD_RETRIES): 7.
  • Max Introduction Scene Retries (INTRO_MAX_RETRIES): 3.
  • Max World Generation Retries (WORLDGEN_MAX_RETRIES): 5.
  • Retry Initial Delay (RETRY_INITIAL_DELAY): 0.25 seconds (for general API call retries).
C.4 Token and API Settings
  • Tokenizer (encoder): cl100k_base (via tiktoken).
  • Max API Token Limit (MAX_API_TOKEN_LIMIT): 60 0006000060\,00060 000 (Safety buffer for Large Language Model calls, allowing space for reasoning tokens if supported by the model endpoint).
  • Max Tokens Buffer (MAX_TOKENS_BUFFER): 500 (Safety margin when checking against MAX_TOTAL_TOKENS).
  • Max Requests Per Second (MAX_REQUESTS_PER_SECOND): 900.0 (Target for OpenRouter rate limiter, dynamically adjusted).
C.5 Generation Cost
[…]
Appendix D Dataset Generation Pipeline
The Verbose ListOps dataset is generated programmatically using an agentic pipeline orchestrated by the verbose-listops.py script. The process for each sample involves:
  1. 1.
    Abstract Syntax Tree Generation: A random ListOps Abstract Syntax Tree (AST) is constructed based on the core ListOps parameters (Appendix C). The Abstract Syntax Tree is then evaluated to determine the ground truth answer.
  2. 2.
    World Generation: An Large Language Model (Gemini 2.5 Flash) generates fictional world metadata (characters, genre, setting, primary object) based on a structured prompt and schema (see Appendix E.1).
  3. 3.
    Narrative Anchor Generation: If USE_NARRATIVE_ANCHORS is true, conceptual names (anchors) for the results of each operation node in the Abstract Syntax Tree are generated, either by an Large Language Model or deterministically.
  4. 4.
    Introduction Scene Generation: An introductory scene is generated by the Large Language Model, setting the stage without revealing numerical details. This scene is validated for numerical compliance (strict zero numbers, with minor exceptions for phrasing).
  5. 5.
    Iterative Beat Generation and Validation:
    • The Abstract Syntax Tree is traversed in post-order. For each OpNode, a narrative "beat" is generated.
    • The Large Language Model generator is provided with a detailed prompt including the current operation, conceptual inputs (anchors from child nodes), new atomic inputs, and an extensive set of "ultra-strict number rules" (see Appendix E.4). These rules enforce that only current atomic operands are stated numerically, prior results are referenced by anchors, and the current operation’s result is implied.
    • The generated beat undergoes an iterative validation loop (_generate_and_llm_validate_beat function):
      1. (a)
        The beat is first validated by another Large Language Model call (LLM_VALIDATOR_MODEL) against the strict rules, using a structured JSON schema for the validator’s response.
      2. (b)
        If the Large Language Model validator fails the beat, its feedback is used to prompt the generator Large Language Model for a revision. This loop continues for up to MAX_LLM_VALIDATION_ITERATIONS.
    • If a beat passes the internal Large Language Model validation loop, it is then subjected to a final Python-based programmatic validation (make_number_validator) to ensure precise numerical compliance (correct numbers mentioned with exact frequencies, no forbidden numbers, result implicitness).
    • If a beat fails either the iterative Large Language Model validation or the final Python validation after all retries (MAX_BEAT_RETRIES for the outer loop), the generation for that entire sample is aborted.
  6. 6.
    Padding Generation: Between valid beats (except after the root node’s beat), optional narrative padding can be inserted to increase context length. Padding is also Large Language Model-generated and validated for numerical compliance (strict zero numbers).
  7. 7.
    Final Question Assembly: A question asking for the final result of the ListOps sequence is appended to the narrative.
  8. 8.
    Output Formatting: Successfully generated samples are saved in JSONL format, including the full narrative, Abstract Syntax Tree, ground truth, and metadata.
The generation process utilizes a ThreadPoolExecutor for parallel generation of multiple samples, with up to DEFAULT_MAX_WORKERS (100 by default). API calls to OpenRouter are managed by a rate limiter.
Appendix E Prompts
This section provides examples of key prompts used in the Verbose ListOps generation pipeline. Note that these are templates and are dynamically filled with specific details (world information, Abstract Syntax Tree node data, etc.) at runtime.
E.1 World Generation Prompt
The Large Language Model is prompted to generate fictional world metadata (characters, genre, setting, object) in a structured JSON format.
System: You are an expert system designed to generate structured data in
**strictly valid JSON format**. Your task is to create fictional world metadata.
**CRITICAL JSON FORMATTING RULES (MUST FOLLOW EXACTLY):**
1.  The entire output MUST be a single, valid JSON object.
2.  All string keys and string values within the JSON must be enclosed in
    double quotes (e.g., "name": "value").
3.  **If a string value itself needs to contain a double quote character
    (e.g., a nickname within a name), that internal double quote MUST be
    escaped with a backslash (‘\\‘)**. For example, if a character’s name is
    ‘Dr. "Nickname" Who‘, it must be represented in the JSON string as
    ‘"name": "Dr. \\"Nickname\\" Who"‘.
4.  Ensure all commas, colons, curly braces ‘{{}}‘, and square brackets ‘[]‘
    are correctly placed according to standard JSON syntax.
5.  Do not include any text, explanations, or markdown (like ‘‘‘json)
    before or after the single JSON object.
**Instructions for Content Generation:**
1.  **Characters:** Generate exactly {num_characters} distinct characters. Each...
    *   ‘name‘: string (e.g., "Kaelen Vane", "Seraphina Moonwhisper")
    *   ‘role‘: string (e.g., "The grizzled warrior," "The cunning sorceress,")
    *   ‘quirk‘: string (e.g., "Collects antique spoons," "Only speaks in riddles,")
2.  **Genre:** Define a ‘genre‘ as a string (e.g., "Steampunk Adventure").
3.  **Setting:** Define a ‘setting‘ as a string (e.g., "A floating city...").
4.  **Object:** Define an ‘object‘ as a string (plural noun, e.g., "etherium crystals").
**Guidance for Content:** Strive for thematic coherence...
Output ONLY the single, valid JSON object.
User: (Dynamically filled with num_characters)
The full prompt includes detailed examples and constraints for each field, ensuring the output adheres to the WORLD_SCHEMA.
E.2 Narrative Anchor Generation Prompt
For OpNodes, conceptual names (anchors) are generated to refer to their results.
System: You are a master {genre} storyteller and creative naming expert.
Your task is to generate a short, evocative, and thematic ’narrative anchor’.
A narrative anchor is a creative, conceptual name that serves as a descriptive
**label** or **stand-in** for the *result* of a specific event or calculation.
Key Guidelines:
1.  **Thematic:** MUST fit Genre, Setting, Primary Object.
2.  **Concise:** 2 to {MAX_ANCHOR_WORDS} words (e.g., ’The Sunstone’s Core’).
3.  **No Numbers:** Absolutely no numerical values.
4.  **No Direct Math Terms:** Avoid ’Sum’, ’Min’, ’Max’, etc.
5.  **Represent Outcome:** Conceptually represent the result.
6.  **Focus on Noun:** Should feel like a "thing" or "state".
7.  **ABSOLUTE UNIQUENESS:** MUST NOT be in ’List of anchors ALREADY USED’.
    If unable, respond with "UNIQUE_FAILURE".
User:
Genre: {genre}
Setting: {setting}
Item: {primary_object}
Concept/Operation Hint: {concept_keywords_for_prompt}
**List of anchors ALREADY USED...:**
{all_previous_anchors}
Provide ONLY the new, unique anchor name... or ’UNIQUE_FAILURE’.
E.3 Introduction Scene Generation Prompt
The introductory scene sets the stage.
System: You are a master {genre} storyteller. Your task is to write a
compelling introductory scene... establish setting, introduce characters,
hint at mystery related to {primary_object}.
**ABSOLUTE NUMERICAL RULE FOR THIS INTRODUCTORY SCENE (CRITICAL):**
1.  **ZERO NUMBERS IS THE PRIMARY GOAL:** Use NO numerical values (digits or words).
2.  **EXTREMELY LIMITED EXCEPTION:** MAY use ’one’, ’two’, or ’three’ for
    general, non-quantitative phrasing IF UNAVOIDABLE. NO OTHER NUMBERS.
3.  **HANDLING CHARACTER NAMES WITH DIGITS:** Avoid stating numerical part as quantity.
    Safer to avoid names with digits for intro.
Focus on atmosphere, intrigue... Output ONLY the narrative text.
User:
**World Context:**
- Genre: {genre}
- Setting: {setting}
- Primary Object of Interest: {object}
- Characters to potentially feature: {char_names_roles}
**Task:** Write an engaging introductory scene...
**CRITICAL REMINDER - ADHERE TO THE ABSOLUTE NUMERICAL RULE...**
Output ONLY the narrative text.
E.4 Main Beat Generation Prompt (Illustrative Core Rules)
This is the most complex prompt, dynamically constructed for each OpNode. The core is the ultra_strict_instruction section. Below is a conceptual summary of its key components:
System: You are a master {genre} storyteller with an exceptional eye for detail...
Your paramount responsibilities for this scene are:
1.  **Narrative Coherence:** ...
2.  **ULTRA-STRICT NUMERICAL AND OPERATIONAL PRECISION:** ...
    *   **Rule 1.A (Exact Atomic Frequencies):** Mention EACH required *new atomic*
        number EXACTLY the specified number of times... AVOID summarizing.
    *   **Rule 1.C (Conceptual Inputs):** Ensure prior results (conceptual inputs)
        are *active numerical inputs* to THIS scene’s operation.
    *   **Operational Fidelity:** Narrated action MUST accurately reflect the
        mathematical operation on ALL inputs.
    *   **Rule 2 (Implicit Outcome):** Numerical result of THIS scene’s operation MUST NOT
        be stated explicitly.
    *   **Rule 4 & 5 (Forbidden & No Other Numbers):** ...
Output ONLY the clean narrative text...
User:
Story Scene Task: Create the narrative for the step resulting in ’{current_node_conceptual_name}’
(Scene {current_beat_num}/{total_beats})
**Background for Your Scene...:**
- Genre: {world_genre}
- Setting: {world_setting}
- Central Items: {primary_object_as_string}
- Quantities from Previous Events (Conceptual Names & values for YOUR understanding -
  DO NOT use values in story, DO use names if Rule 1.C applies): {conceptual_inputs_context_str}
- New Numbers Introduced (Values & required frequencies for YOUR understanding - Use word
  form, ALL must be mentioned with exact frequencies as per Rule 1.A): {atomic_inputs_context_str_detailed_for_prompt}
**Your Scene’s Core Action & Narrative Goal (Follow this closely):**
This scene needs to narrate an event or discovery that mirrors the mathematical
operation: **{op_label}**. The central items are ’{safe_primary_object_for_fstring}’.
Inputs to consider:
  1. **Conceptual Inputs:** {conceptual_input_names_only_str_for_action}.
  2. **New Atomic Number Inputs:** {atomic_inputs_context_str_detailed_for_prompt}.
Your narrative must clearly show ALL these inputs... being involved in an action
that reflects the ’{op_label}’ operation.
  - **Action (Specific to Op, e.g., SUM):** Characters combine/tally ALL inputs...
The outcome will be conceptually known as ’{current_node_conceptual_name}’.
Its actual numerical size (’{num_to_words(correct_result)}’) must not be stated.
**Narrative Challenge & Your Writing Guide for This Scene (CRITICAL...):**
**1. Key Details to Feature (Inputs in Action & Their EXACT Frequencies):**
[...refer to codebase for full prompt...]
{prior_results_handling_rule_for_prompt} (This is Rule 6)
**Operational Fidelity (CRITICAL):** The narrated action MUST accurately reflect
the mathematical operation ’{op_label}’ on ALL inputs (conceptual & atomic).
**MANDATORY PRE-WRITING CHECKLIST & MENTAL WALKTHROUGH...:**
(Detailed checklist items for the LLM to mentally verify its plan against each rule)
**Continue From (End of last scene):**
"...{context_snippet}..."
**Your Response:**
Write ONLY the narrative text for this new scene...
The actual prompt is highly detailed, including specific examples for MEDIAN operations and a pre-writing checklist for the Large Language Model. The ultra_strict_instruction section is dynamically built based on the current node’s operation, its inputs (atomic and conceptual), its result, and the overall Abstract Syntax Tree context to define precisely which numbers are allowed, required (with exact frequencies), or forbidden for that specific beat.
E.5 Large Language Model Validator Prompt (Iterative Beat Validation)
During the iterative beat generation, another Large Language Model validates the generator’s output.
System: You are an AI numerical compliance checker and literary critic.
Your ONLY task is to evaluate a story ’beat’ against a provided set of
ULTRA-STRICT numerical and storytelling rules.
You MUST output your response as a single, valid JSON object and NOTHING ELSE,
adhering precisely to the provided schema.
Your analysis must be meticulous, focusing on exact numerical frequencies...
User:
You are an AI numerical compliance checker. Evaluate the ’Generated Beat Text’
below with ABSOLUTE PRECISION regarding its numerical content, operational
fidelity, and adherence to the ’ULTRA-STRICT NUMBER RULES (Generator’s
Writing Guide)’ provided.
[...see codebase for full prompt...]
**ULTRA-STRICT NUMBER RULES (Generator’s Writing Guide - GROUND TRUTH FOR VALIDATION):**
{ultra_strict_instruction_for_llm_validator_context} (This is the full ruleset given to the generator)
**VALIDATION ALGORITHM - FOLLOW EXACTLY...:**
(Detailed step-by-step algorithm for the validator LLM to check each rule:
 Phase 1: Number Identification & Counting.
 Phase 2: Rule-by-Rule Compliance Check (Rule 0.A Conceptual Inputs,
          0.C Operational Fidelity, 1.A Atomic Frequencies, 1.B No Re-listing,
          2 Outcome Handling, 3 Permitted Flourishes, 4 Forbidden, 5 No Other,
          6 Prior Result Handling).
 Phase 3: Constructing JSON Response according to VALIDATOR_RESPONSE_SCHEMA,
          including ‘is_valid‘, ‘explanation_for_generator‘, ‘explanation_for_audit‘,
          ‘overall_revision_summary_for_generator_prompt‘, ‘suggested_revisions‘.)
**Generated Beat Text to Evaluate:**
{generated_text_cleaned}
The validator’s prompt includes the full set of rules given to the generator, the generated text, and a detailed algorithm for how the validator should check compliance and structure its JSON response.
E.6 Final Question Template
The template for the final question appended to the narrative.
FINAL_QUESTION_TEMPLATE = Template(
[…]
    "modify a quantifiable measure related to ’\$primary_object’. Following this "
    "entire sequence, what is the final, precise numerical value of this measure "
    "at the conclusion of all activities? Provide only the single integer."
Here, \$primary_object is substituted with the specific object generated for the world (e.g., "etherium crystals").
Appendix F Dataset Details and Access
The Verbose ListOps dataset is available on Hugging Face Datasets:
The dataset is provided in JSON Lines (.jsonl) format.
  • [1_RESEARCHER_DETAIL]_DATASET_.jsonl: Contains comprehensive raw generated data for each sample, including detailed generation metadata, full Abstract Syntax Tree structure, all scenes, conceptual references, and beat revision logs. This file is primarily for research and debugging the generation process.
  • [2_EVAL_READY]_DATASET_.jsonl: A leaner version containing all successfully generated samples by verbose-listops.py, with fields relevant for model evaluation. This is the dataset before external validation by validator.py.
  • [4_FINAL_EVAL_CLEANED]_DATASET_.jsonl: The final, cleaned dataset intended for benchmarking. This version contains only samples that have passed an additional validation step by the validator.py script, ensuring higher fidelity of the narrative to the underlying ListOps task according to an external Large Language Model judge.
[…]
F.1 Key Dataset Fields (in EVAL_READY and FINAL_EVAL_CLEANED versions)
Table 4 describes the main fields in the evaluation-ready JSONL files.
Table 4: Key fields in the Verbose ListOps evaluation-ready dataset files.
Field Name Description
id Unique identifier for the sample (string).
full_text_for_eval The complete text provided to the Large Language Model for evaluation, consisting of the narrative body followed by the final question (string).
ground_truth_value The single integer answer to the ListOps problem (integer).
ast_str A string representation of the ListOps Abstract Syntax Tree in prefix notation (string).
num_operations The total number of operation nodes (non-atom nodes) in the Abstract Syntax Tree (integer).
token_count_narrative The approximate token count of the narrative body (excluding the question), based on cl100k_base tokenizer (integer).
The RESEARCHER_DETAIL file contains additional fields like world_data, scenes_detail, conceptual_references, beat_revision_details, and extensive generation_metadata.
Appendix G Evaluation Details
  • Evaluation Metric: Performance is measured by exact match accuracy. The Large Language Model’s predicted single integer answer must exactly match the ground_truth_value for the sample.
[…]
  • External Validation (validator.py): The validator.py script performs an additional layer of validation on the generated narratives. It uses a separate, more powerful thinking Large Language Model (google/gemini-2.5-pro-preview as specified in validator.py) with extensive prompting to assess whether each step in the narrative correctly reflects the corresponding Abstract Syntax Tree operation, its inputs, and implied result, according to the benchmark’s rules (including implicit intermediate results and conceptual referencing). Samples that fail this external validation are excluded from the FINAL_EVAL_CLEANED dataset. This script outputs detailed validation results and helps ensure the quality and fidelity of the benchmark instances used for final model evaluation.
Appendix H Code Availability
The code for generating the Verbose ListOps benchmark and the validator.py script are open-sourced and available at:
The repository contains the verbose-listops.py script for dataset generation and the validator.py script for post-generation validation and cleaning. Detailed instructions for running the scripts and reproducing the dataset are provided in the repository’s README file.

### v4

@@
 Abstract
-The dominant evaluation of Large Language Models has centered on their ability to surface explicit facts from increasingly vast contexts.
-While today’s best models demonstrate near-perfect recall on these tasks, this apparent success masks a fundamental failure in multi-step computation when information is embedded in a narrative.
-We introduce Verbose ListOps (VLO), a novel benchmark designed to isolate this failure.
-VLO programmatically weaves deterministic, nested computations into coherent stories, forcing models to track and update internal state rather than simply locate explicit values.
-Our experiments show that leading LLMs, capable of solving the raw ListOps equations with near-perfect accuracy, collapse in performance on VLO at just 10k tokens.
-The VLO framework is extensible to any verifiable reasoning task, providing a critical tool to move beyond simply expanding context windows and begin building models with the robust, stateful comprehension required for complex knowledge work.
+The dominant way of judging Large Language Models (LLMs) has been to ask how well they can recall explicit facts from very long inputs.
+While today’s best models achieve near-perfect recall, this masks a harder skill: performing multi-step reasoning and tracking intermediate state that never appears verbatim.
+We introduce Verbose ListOps (VLO), a benchmark that embeds deterministic ListOps computations inside narrative camouflage and, crucially, allows step-level evaluation of every intermediate result.
+Experiments show that models which solve raw ListOps with $\approx 100\%$accuracy collapse on VLO after only 10 k tokens.
+By exposing where a model’s reasoning chain first diverges, VLO moves assessment beyond sheer context length and toward genuine comprehension.
+VLO’s generation pipeline is task-agnostic: it can weave any deterministically verifiable reasoning schema—arithmetic, symbolic, abductive, inductive or defeasible—into narrative form.
+This makes VLO a reusable test-bed for the next wave of reasoning-centric model designs, not merely those with step-explicit scaffolds.
 1 Introduction
-Despite boasting million-token context windows, today’s leading Large Language Models (LLMs) can fail at reasoning tasks a human finds trivial: tracking a multi-step argument buried within a distracting narrative.
-This ability—to filter irrelevant information and track intermediate conclusions for later synthesis—is the foundation of high-value knowledge work, from a lawyer interpreting intepreting clauses to a sales manager inferring customer intent from chatty transcripts.
-Yet this core skill remains a fundamental limitation as existing benchmarks are ill-equipped to measure it.
+Despite million-token context windows, today’s LLMs still stumble on tasks that a human finds trivial: following a multi-step argument buried inside a distracting story (Liu et al., 2024; Shi et al., 2023; An et al., 2023).
+This ability—to ignore irrelevant detail, maintain intermediate conclusions, and synthesise them later—is the backbone of knowledge work, from a lawyer interpreting clauses to a sales manager inferring intent from chatty transcripts.
+Yet benchmarks rarely measure it.
 To bridge this gap, we introduce Verbose ListOps (VLO), a novel benchmark that isolates this ability in a controlled setting, challenging models to reason, not just locate information.
+Existing long-context suites chiefly test (i) factual recall in "needle-in-a-haystack" settings (Kamradt, 2023; Li et al., 2024b) or (ii) shallow reading comprehension (An et al., 2023; Zhang et al., 2024).
+They do not control reasoning depth, nor do they hide intermediate answers, so models can succeed through pattern matching rather than computation.
+Verbose ListOps (VLO) fills this gap.
+ListOps is a synthetic benchmark where models perform nested arithmetic operations on sequences of integers following a tree structure—for example, [MAX 3 7 [MIN 1 9 5] 2] requires computing the inner MIN first, then finding the maximum among all values.
+We convert each ListOps node into a short story fragment, hide every intermediate value behind a ’narrative anchor’, and add semantically related distractors.
+The resulting benchmark forces a model to compute, cache, and reuse internal state.
 Figure 1: VLO embeds ListOps problems in a narrative, forcing LLMs to find and solve each subproblem, track intermediate states, and ignore distractors for the final answer, emulating complex human text synthesis.
 Takeaway: VLO reveals LLMs needs more than just long-context processing—they must compute within distracting narratives where intermediate results remain implicit.
-VLO is specifically designed to address the limitations of prior evaluation methods.
-Where benchmarks for sequence comprehension (An et al., 2023; Zhang et al., 2024) or needle-in-a-haystack factual recall (Kamradt, 2023; Li et al., 2024b) test factual extraction, VLO embeds deterministic, nested ListOps computations (Nangia and Bowman, 2018; Tay et al., 2021) inside lengthy, coherent stories.
-Where synthetic datasets can leak intermediate solutions that models merely recall (Fodor, 2025), VLO enforces a strict protocol that withholds every intermediate result, forcing models to compute and maintain values in ’working memory’.
-And while human-annotated datasets offer realism, they are too coarse-grained to disentangle narrative interference from reasoning difficulty (Wang et al., 2024; Shaham et al., 2022; Bowman and Dahl, 2021); VLO’s agentic generation provides orthogonal controls for both.
-This setup probes an LLM’s ability to maintain computational ’state’ and follow algorithmic logic under narrative camouflage—coherent, semantically-related distractors that obscure a task—prerequisites for text understanding comparable to human performance.
+In this paper, we deliver three key contributions:
+  1. 1.
+    A broadly-applicable, agentic generation framework.
+We open-source a "World-Author–Critic–Validator" pipeline that can embed any deterministically-verifiable reasoning graph—not only arithmetic ListOps but also symbolic, abductive, inductive, or defeasible logic (see Appendix A).
+  2. 2.
+    Verbose ListOps (VLO): a first instantiation of the framework.
+Using this pipeline we build VLO, a benchmark that narrativises nested ListOps computations.
+VLO is the first long-context dataset that (i) withholds all intermediate results, (ii) controls context length and reasoning depth orthogonally, and (iii) ships exact intermediate-step labels, enabling systematic failure-mode analysis.
+  3. 3.
+    Orthogonal control knobs for context and reasoning complexity.
+Tree depth, operator arity, narrative length, and distraction density are exposed as independent hyper-parameters, allowing researchers to stress-test scaling behaviour along a single axis while holding all others fixed—something prior benchmark suites cannot do.
+  4. 4.
+    A diagnostic evaluation toolkit and empirical study.
+We release scripts that check correctness at every anchor step, plot error-propagation curves, and compare prompting regimes.
+Applying this toolkit to ten prominent LLMs uncovers a consistent blind spot: models that excel at raw ListOps or factual recall often diverge at the very first hidden step when the same logic is wrapped in narrative camouflage.
+Together, the framework plus toolkit turn VLO from a one-off dataset into a reusable test-bed for any future long-context reasoning architecture.
+Our experiments show state-of-the-art LLMs, despite solving ListOps with ease, on VLO, suffer a severe ($\approx$50%+) drop in accuracy at just $\approx$10k-token contexts.
+The result exposes an under-tested limitation: LLMs struggle to synthesize multi-step conclusions amongst narrative distraction.
+VLO offers a rigorous, open-source tool for diagnosing (and improving) this foundational capability, seeking to move discussion beyond simply enlarging context windows.
+2 Related Work
+The expansion of LLM context windows has catalyzed research into their performance on extensive textual inputs, leading to a diverse landscape of evaluation benchmarks.
 Benchmark Primary Reasoning Task Nested Reasoning Chains Distraction Tunable Reasoning Difficulty Scalable Context Deterministic Answer Realistic Tasks Generation Method
@@
 Table 1: Comparison of Verbose ListOps with other long-context benchmarks. Using an agentic generation process, Verbose ListOps offers both controllable context lengths and reasoning difficulty.
-In this paper, we deliver three key contributions:
-  1. 1.
-    A benchmark for narrative-embedded deterministic reasoning.
-We weave ListOps computations into coherent fictional stories and withhold the intermediate results of all the nested problems, forcing models to compute, store, and recall values internally.
-  2. 2.
-    Orthogonal control of context length and reasoning complexity.
-We provide independent parameters for narrative length and reasoning complexity, enabling systemic exploration of scaling behavior and failure modes.
-  3. 3.
-    An extensible generalizable generation framework.
-We open-source a fully programmatic agentic pipeline where "author" and "critic" LLMs collaborate to embed any deterministically verifiable task; from numerical operations to symbolic and logical reasoning, into a coherent narrative.
-This methodology yields reliable, scalable data and provides a foundation for testing a wide range of complex reasoning beyond the scope of this initial work.
-Our experiments show state-of-the-art LLMs, despite solving ListOps with ease, on VLO, suffer a severe ($\approx$50%+) drop in accuracy at just $\approx$10k-token contexts.
-The result exposes an under-tested limitation: LLMs struggle to synthesize multi-step conclusions amongst narrative distraction.
-VLO offers a rigorous, open-source tool for diagnosing (and improving) this foundational capability, seeking to move discussion beyond simply enlarging context windows.
-2 Related Work
-The expansion of LLM context windows has catalyzed research into their performance on extensive textual inputs, leading to a diverse landscape of evaluation benchmarks.
 2.1 Long-Context Benchmarks
 Early long-context evaluations typically adapted standard NLP benchmarks (Shaham et al., 2023; An et al., 2023; Bai et al., 2024).
-These early benchmarks often had contexts shorter than the maximum capabilities of current LLMs (Gemini Team, 2024; OpenAI, 2025a; Anthropic, 2025) and inadequately differentiated complex reasoning tasks from simpler simpler tasks based on factual recall tasks or failed to analyze the impact of distractors sufficiently (Muhlgay et al., 2023).
+These early benchmarks often had contexts shorter than the maximum capabilities of current LLMs (Gemini Team, 2024; OpenAI, 2025a; Anthropic, 2025) and inadequately differentiated complex reasoning tasks from simpler tasks based on factual recall tasks or failed to analyze the impact of distractors sufficiently (Muhlgay et al., 2023).
 Recent synthetic benchmarks have addressed some of these limitations by allowing greater control over context length.
@@
 Likewise, the LaRA benchmark (2025) finds long-context LLMs outperform RAG on multi-step arithmetic problems as document length scales (Li et al., 2025).
+As VLO exposes every intermediate state, it is also a natural playground for Process-Reward Models (PRMs) that score reasoning chains rather than final outputs.
+We expect the dataset to serve both as a training signal and as a forensic tool for such chain-inspection methods.
 3 Verbose ListOps (VLO): Specification and Construction
@@
 For example, if
-$T=\textsc{max}(\textsc{sum}(2,1),\,4),$
-then one computes $\textsc{sum}(2,1)=3$ and subsequently $\textsc{max}(3,4)=4$.
+$T=\texttt{MAX}(\texttt{SUM}(2,1),\,4),$
+then one computes $\texttt{SUM}(2,1)=3$ and subsequently $\texttt{MAX}(3,4)=4$.
 VLO embeds each nested ListOps operation in a narrative, describing each node in $T$ via post-order traversal.
@@
 $f\colon X\longrightarrow v(T),$
-with intermediate values appearing only as anchors ${a_{i}}$. Success is $f(X)$ matching the ground-truth integer.
+…where v(T) denotes the scalar value at the root of the abstract-syntax tree (i.e. the final numerical result). Success is $f(X)$ matching the ground-truth integer.
 Running example
-Consider the above ListOps problem again. The first operation is
-$\textsc{sum}(2,1)\;\to\;a_{1},\quad a_{1}=3$
-This intermediate value (3) is tracked by anchor $a_{1}$. Example narrative segment:
+Consider the above ListOps problem again. Intermediate values appear only as anchors ${a_{i}}$. The first operation is
+$\texttt{SUM}(2,1)\;\to\;a_{1},\quad a_{1}=3$
+In VLO, this operation would be narrativized as the following:
 "In the bustling spaceport of Xylos, Chief Engineer Anya cataloged incoming parts.
@@
 The second operation is
-$\textsc{max}(a_{1},4)\;\to\;a_{2},\quad a_{2}=4$
+$\texttt{MAX}(a_{1},4)\;\to\;a_{2},\quad a_{2}=4$
 …yielding the final result $v(T)$.
@@
 3.3 Generation Pipeline
-Figure 2 outlines how the VLO construction pipeline ensures adherence to constraints C1–C3 through:
+Figure 2 outlines how the VLO construction pipeline ensures adherence to constraints C1–C3 through the following five-stage process (implementation details and prompts are provided in Appendices D and E):
   1. 1.
-    AST Sampling (Programmatic). Sample a ListOps AST $T$ of depth $d$ and branching factor $b$, and compute its root value $v(T)$ by deterministic post-order evaluation.
+    AST Sampling. Sample a ListOps AST $T$ of depth $d$ and branching factor $b$, and compute its root value $v(T)$ by deterministic post-order evaluation.
   2. 2.
-    Iterative Operator Narrative Construction. For each ListOps node, build a narrative segment via:
+    World Generation.
+Generate fictional world metadata (genre, setting, characters, and primary objects) using an LLM to establish thematic coherence for narrative embedding (constraint C2).
+This creates the contextual foundation for embedding ListOps operations into a cohesive story.
+Narrative anchors—conceptual names like "Daily Cell Intake"—are generated to reference intermediate results without revealing numerical values (constraint C1).
+Detailed parameters are in Appendix C.
+  3. 3.
+    Distraction Generation.
+Generate semantically relevant but computation-irrelevant content between ListOps operations.
+This expands context length while forcing models to filter computational information from extraneous narrative detail, using the same world context to maintain coherence without introducing interfering numerical information that can interfere with the underlying reasoning task.
+The token watcher monitors narrative length throughout generation, dynamically allocating the remaining token budget after ListOps narratives.
+  4. 4.
+    Operator Narrative Construction. For each ListOps node, build a narrative segment via:
     • (a)
-      Author Agent (LLM).
+      Author Agent.
 The Author LLM drafts a segment for the current operator $\omega\in\mathcal{O}$ and its child anchors $a_{i}$ (if applicable), weaving it into the story (C2) while avoiding numeric mention of the current result (C1) and respecting parameters (C3).
     • (b)
-      Critic Agent (LLM). The Critic LLM reviews the draft, checking:
+      Critic Agent. The Critic LLM reviews the draft, checking:
       • No numeric digits or words for the current operation’s result appear (enforces C1).
@@
     • (c)
-      Static Validation (Programmatic). A script scans the finalized segment to verify:
+      Programmatic Static Validation. A script scans the finalized segment to verify:
       • Absence of any digit or word numbers for the current operation’s result.
@@
       Any failure triggers a re-generation of that segment.
-  3. 3.
-    Holistic Narrative Validation (LLM). After concatenating all segments into the full narrative, validator.py orchestrates a review of the entire sample:
-    • Comprehensive LLM validation: A larger state-of-the-art Validator LLM, aided by extensive prompting and few-shot examples, reads the entire story.
-It performs a step-by-step evaluation of the narrative against the original AST $T$ to detect subtle numeric leaks (e.g., spelled-out numbers for intermediate results), verify correct representation of operations and inputs (including conceptual anchors), check for missing or duplicate operators, ensure overall coherence (C2), and confirm that its own evaluation of the original AST $T$ matches the provided ground truth value $v(T)$.
+  5. 5.
+    Comprehensive LLM Validation: A larger Validator LLM, aided by extensive prompting and few-shot examples, performs step-by-step evaluation against the original AST $T$ to detect numeric leaks, verify correct representation of operations and inputs, check for missing/duplicate operators, ensure coherence (C2), and confirm ground truth value $v(T)$.
 Violations identified by the Validator LLM cause the sample to be discarded and resampled.
-Figure 2: The Verbose ListOps (VLO) agentic generation and validation pipeline, employing Author, Critic, and Validator LLMs alongside programmatic checks to ensure sample validity.
+Figure 2: The Verbose ListOps (VLO) agentic generation and validation pipeline.
+The process begins with AST sampling and world generation, then iteratively constructs narrative segments using Author-Critic LLM pairs with programmatic validation, incorporates distraction content, and concludes with holistic validation to ensure sample validity and adherence to constraints C1-C3.
 3.4 Controllable Parameters
@@
 Whilst nearly every ’thinking’ model achieves perfect or near-perfect accuracy on bare ListOps, performance collapses under Verbose ListOps at just 10-k tokens.
-Model ListOps VLO-10k OAI MRCR-8k 8-Pin 95% CI (ListOps) 95% CI (VLO)
+Model ListOps VLO-10k OAI MRCR-8k 8-Pin [^1] OpenAI’s Multi-Reference Context Recall (MRCR) benchmark: 8-passage subset measuring pure retrieval accuracy under 8k tokens without multi-step reasoning.
+Provides baseline for comparing retrieval vs.
+VLO’s narrative-embedded reasoning.
+Results from ContextArena (Uzar, 2025).
+95% CI (ListOps) 95% CI (VLO)
 Closed-source models
@@
 DeepSeek V3 0324 93.7 25.1 N/A 92.0–95.1 22.4–28.0
-Table 2: Accuracy on VLO-10k versus its corresponding bare ListOps expressions, based on evaluation of 1,000 samples per model.
-Of particular interest is DeepSeek V3’s plunge in performance from 93.7% to 25.1%.
-OpenAI MRCR results (95% CI) sourced from ContextArena (Uzar, 2025).
+Table 2: Accuracy on VLO-10k versus its corresponding bare ListOps expressions, based on evaluation of 1,000 samples per model. Note DeepSeek V3’s drop from 93.7% to 25.1%.
 3.8 Analysis and Limitations
 Narrative Camouflage, ’Cognitive Control’, and Architectural Divergence
-The VLO benchmark’s core challenge is its use of narrative camouflage to obscure a nested deterministic task, creating a information processing burden that exposes fundamental architectural differences between models.
+VLO challenges models to find, solve, track, and synthesize nested reasoning problems amongst narrative camouflage to answer a deterministic question, creating a information processing burden that exposes fundamental architectural differences between models.
 The results in Table 2 reveal two critical takeaways.
@@
 DeepSeek-V3’s design, while highly efficient, creates specific vulnerabilities to VLO’s challenge.
-Its architecture is optimized for performance on standard benchmarks through two key strategies that, we hypothesize, compromise its task resiliance on VLO.
+Its architecture is optimized for performance on standard benchmarks through two key strategies that, we hypothesize, compromise its task resilience on VLO.
 First, its auxiliary-loss-free load balancing for its Mixture-of-Experts (MoE) architecture encourages over-specialisation, where the gating network routes inputs to a "narrative" expert that fails to pattern-match the embedded logic.
@@
 Further considerations and avenues for future research include:
-  • Generator Model Bias: VLO-10k was generated with Gemini 2.5.111During development we found only Gemini 2.5 and GPT-4.5 could reliably generate VLO, with GPT-4.5 pricing being cost-prohibitive (\$75USD/1M input tokens, \$150USD/1M output tokens (OpenAI, 2025b)).
+  • Generator Model Bias: VLO-10k was generated with Gemini 2.5.222During development we found only Gemini 2.5 and GPT-4.5 could reliably generate VLO, with GPT-4.5 pricing being cost-prohibitive (\$75USD/M input tokens, \$150USD/M output tokens (OpenAI, 2025b)).
 This introduces potential generator model bias, where evaluated models from the same family may exhibit inflated performance due to stylistic or implicit knowledge alignment.
@@
 (2023)’s Tree-of-Thoughts.
-Concept-oriented Large Concept Models (LCMs),222’LCMs’ here do not refer to Latent Consistency Models (Luo et al., 2023) used for image synthesis.
+Concept-oriented Large Concept Models (LCMs),333’LCMs’ here do not refer to Latent Consistency Models (Luo et al., 2023) used for image synthesis.
 which operate on higher-level semantic units rather than individual tokens (Meta LCM Team et al., 2024) offer a promising direction.
@@
   • Example AST:
-    (ABDUCE
+[…]
       (OBSERVATIONS "lights flicker" "strange hum")
@@
   • Example AST:
-    (INDUCE
+[…]
       (EXAMPLES ("raven A is black") ("raven B is black"))
@@
   • Example AST:
-    (DEFEASIBLE_QUERY
+[…]
       (QUERY "Tweety can fly")
@@
 3.  **No Numbers:** Absolutely no numerical values.
-4.  **No Direct Math Terms:** Avoid ’Sum’, ’Min’, ’Max’, etc.
+[…]
 5.  **Represent Outcome:** Conceptually represent the result.
@@
 that reflects the ’{op_label}’ operation.
-  - **Action (Specific to Op, e.g., SUM):** Characters combine/tally ALL inputs...
+[…]
 The outcome will be conceptually known as ’{current_node_conceptual_name}’.